find the inverse of this matrix


A= 1 3 5 4


7 2 9 6


1
Expert's answer
2022-01-18T10:42:53-0500
"A=\begin{bmatrix}
 1 & 3 & 5 \\
 4 & 7 & 2 \\
9 & 6 & 8
\end{bmatrix}"

"\begin{bmatrix}
 1 & 3 & 5 & & 1 & 0 & 0 \\
 4 & 7 & 2 & & 0 & 1 & 0 \\
 9 & 6 & 8 & & 0 & 0 & 1 
\end{bmatrix}"

"R_2=R_2-4R_1"


"\begin{bmatrix}
 1 & 3 & 5 & & 1 & 0 & 0 \\
 0 & -5 & -18 & & -4 & 1 & 0 \\
 9 & 6 & 8 & & 0 & 0 & 1 
\end{bmatrix}"

"R_3=R_3-9R_1"


"\begin{bmatrix}
 1 & 3 & 5 & & 1 & 0 & 0 \\
 0 & -5 & -18 & & -4 & 1 & 0 \\
 0 & -21 & -37 & & -9 & 0 & 1 
\end{bmatrix}"

"R_2=-R_2\/5"


"\begin{bmatrix}
 1 & 3 & 5 & & 1 & 0 & 0 \\
 0 & 1 & 18/5 & & 4/5 & -1/5 & 0 \\
 0 & -21 & -37 & & -9 & 0 & 1 
\end{bmatrix}"

"R_1=R_1-3R_2"


"\begin{bmatrix}
 1 & 0 & -29/5 & & -7/5 & 3/5 & 0 \\
 0 & 1 & 18/5 & & 4/5 & -1/5 & 0 \\
 0 & -21 & -37 & & -9 & 0 & 1 
\end{bmatrix}"

"R_3=R_3+21R_2"


"\begin{bmatrix}
 1 & 0 & -29/5 & & -7/5 & 3/5 & 0 \\
 0 & 1 & 18/5 & & 4/5 & -1/5 & 0 \\
 0 & 0 & 193/5 & & 39/5 & -21/5 & 1 
\end{bmatrix}"

"R_3=5R_3\/193"


"\begin{bmatrix}
 1 & 0 & -29/5 & & -7/5 & 3/5 & 0 \\
 0 & 1 & 18/5 & & 4/5 & -1/5 & 0 \\
 0 & 0 & 1 & & 39/193 & -21/193 & 5/193 
\end{bmatrix}"

"R_1=R_1+29R_3\/5"


"\begin{bmatrix}
 1 & 0 & 0 & & -44/193 & -6/193 & 29/193 \\
 0 & 1 & 18/5 & & 4/5 & -1/5 & 0 \\
 0 & 0 & 1 & & 39/193 & -21/193 & 5/193 
\end{bmatrix}"

"R_2=R_2-18R_3\/5"


"\begin{bmatrix}
 1 & 0 & 0 & & -44/193 & -6/193 & 29/193 \\
 0 & 1 & 0 & & 14/193 & 37/193 & -18/193 \\
 0 & 0 & 1 & & 39/193 & -21/193 & 5/193 
\end{bmatrix}"

We are done. On the left is the identity matrix. On the right is the inverse matrix.


"A^{-1}=\begin{bmatrix}
 -44/193 & -6/193 & 29/193 \\
 14/193 & 37/193 & -18/193 \\
39/193 & -21/193 & 5/193
\end{bmatrix}"